### v2

# Neural Measures for Learning Distributions of Random PDEs

Abstract
The integration of Scientific Machine Learning (SciML) techniques with uncertainty quantification (UQ) represents a rapidly evolving frontier in computational science. This work advances Physics-Informed Neural Networks (PINNs) by incorporating probabilistic frameworks to effectively model uncertainty in complex systems. Our approach enhances the representation of uncertainty in forward problems by combining generative modeling techniques with PINNs. This integration enables in a systematic fashion uncertainty control while maintaining the predictive accuracy of the model. We demonstrate the utility of this method through applications to random differential equations and random partial differential equations (PDEs).
keywords:
Random PDEs, Neural networks for PDEs, Generative models, Neural measure, Physics-Informed Neural Networks, Universal Approximation Theorem, Uncertainty Quantification.
{MSCcodes}
65C20, 60H35, 35R60
1 Introduction
1.1 Problem Formulation and Approach
In this paper, we consider random differential equations, which are differential equations that depend on random parameters. Assuming that we have knowledge of the probability distribution of these parameters, we aim to assess the probability distribution of the solutions. This is a fundamental problem in uncertainty quantification and can be regarded as the initial step in designing algorithms for more intricate inverse problems where we aim to determine both the form of the equation and its solutions with controlled uncertainty by leveraging the available data.
The primary tools we will employ include discrete neural network methods and generative models, along with adaptations of ideas from Physics-Informed Neural Networks tailored to our current problem. Machine learning and neural networks have proven to be highly effective tools for approximating probability distributions in $\mathbb{R}^{k}$, even for large dimensions $k$. However, the solutions of random differential equations can be viewed as probability distributions on infinite-dimensional spaces (typically Hilbert or Sobolev spaces), making their formulation and approximation particularly intricate. This is one of the key issues we aim to address in the present work.
The problems we consider have the form
(1) $\begin{split}&\mathscr{A}_{\xi}u(\cdot,\xi)=\mathscr{F}_{\xi},\end{split}$
where $\mathscr{A}_{\xi}$ is a $\xi$-parametrised differential operator (incorporating possible initial and boundary conditions), and $\mathscr{F}_{\xi}$ is the source. For each fixed $\xi$, $u(\cdot;\xi)$ denotes the solution of the differential equation, which we interpret as a mapping $U:\Xi\to\mathcal{H}$ with $U(\xi)\coloneqq u(\cdot,\xi)$. The space $\mathcal{H}$ where $u$ takes values for each $\xi$ is an (infinite dimensional) Hilbert or Sobolev space. We assume that the random parameter $\xi\in\Xi$ is distributed according to a known Borel probability measure $\gamma$ on $\Xi$. Our aim is to approximate the random variable $U$ and its law $\nu_{u}\in\mathcal{P}(\mathcal{H})\,;$ here $\mathcal{P}(\mathscr{X})$ denotes the set of Borel probability measures on $\mathscr{X}$.
Motivation and Neural Measures
One of the key successes of neural network methods is their ability to effectively approximate probability distributions. In fact, several important problems of interest rely on generating a probability distribution when the available data represent specific instances of the unknown distribution. Typically, if the data is represented by an empirical measure $\nu_{\,\text{data}}\in\mathcal{P}(\mathbb{R}^{k})$, we would like to generate an approximation of $\nu_{\,\text{true}}\in\mathcal{P}(\mathbb{R}^{k})$ by optimising
(2) $\min_{\mu\in V_{\Theta}}d_{\mathscr{M}}\left(\mu,\nu_{\,\text{data}}\right),$
where $V_{\Theta}$ are appropriate discrete subsets of $\mathcal{P}(\mathbb{R}^{k})$ determined by a parameter set $\Theta$ which is connected in an appropriate manner with neural network sets of given architecture and $d_{\mathscr{M}}\left(\cdot,\cdot\right)$ a specified distance (or divergence) on $\mathcal{P}(\mathbb{R}^{k})$. We then hope that
(3) $\mu_{\theta^{*}}\cong\nu_{\,\text{true}},\qquad\text{where }\mu_{\theta^{*}}\in\operatorname*{arg\,min}_{\mu\in V_{\Theta}}d_{\mathscr{M}}\left(\mu,\nu_{\,\text{data}}\right).$
Representative works utilising the aforementioned generative model framework for a variety of applications include [1, 17, 4, 59, 9, 19], and the references in [55]. For a recent modelling perspective that interprets various generative models as mean field games, see [55]. Recent advances connect random PDEs with foundational models [56].
The above framework relies on various alternative choices, each one yielding fundamentally distinct algorithms at the end. We would like to explore the potential extension of this approach to approximate solutions of Eq. 1. To accomplish this we must first systematically construct discrete spaces $V_{\Theta}$ such that $V_{\Theta}\subset\mathcal{P}(\mathcal{H}).$ We adopt the following alternative approaches:
  • fully NN based $V_{\Theta},$ detailed in Section 3.1
  • Polynomial Chaos Expansion (PCE)-NN $V_{\Theta},$ detailed in Section 3.2
  • Galerkin-NN $V_{\Theta},$ detailed in Section 3.3.
Furthermore, we show that the neural measure spaces are expressive enough to approximate appropriate subspaces of $\mathcal{P}(\mathcal{H})$.
Incorporating PINNs
Once appropriate $V_{\Theta},$ $V_{\Theta}\subset\mathcal{P}(\mathcal{H}),$ are in place, one may want to incorporate the key idea of Physics Informed Neural Networks and to combine it with the Neural Measure spaces. In fact, one needs to define
(4) $\mathscr{A}\odot\mu,\qquad\mu\in V_{\Theta}\subset\mathcal{P}(\mathcal{H}),$
i.e. the probability distribution on $\mathcal{P}(\mathcal{H}^{\text{imag}}),$ which is obtained by applying $\mathscr{A}_{\xi}$ to $\mu$ in an appropriate sense. Here $\mathscr{A}_{\xi}u(\cdot,\xi),\mathscr{F}_{\xi}\in\mathcal{H}^{\text{imag}}$ for each fixed $\xi.$ This definition is made precise in Sections 2.3 and 2.4.
We then consider the abstract problem,
(5) $\min_{\mu\in V_{\Theta}}d_{\tilde{\mathscr{M}}}\left(\mathscr{A}\odot\mu,\nu_{\,\mathscr{F}}\right),$
[…]
Formulation of the methods and results
The plan outlined above is implemented systematically in the following sections. We adopt a concrete yet somewhat abstract approach for two main reasons:
(a) the formulation of discrete neural measure spaces has the potential to be valuable in various other applications of uncertainty quantification (UQ) for partial differential equations (PDEs), and
(b) the integration of the physics-informed neural networks (PINN) framework into the final algorithms can be carried out in a general manner, independent of the specific choice of measure metrics.
We then show that employing Wasserstein distances offers particular advantages in the implementation of the final algorithms.
This work is structured to systematically develop and analyse a generative modelling approach for Uncertainty Quantification using Physics-Informed Neural Networks (PINNs). In Section 2, we introduce the underlying problem, and the corresponding probabilistic framework essential to our study. Next, we introduce new notational conventions and the variational formulation used. The variational principle forms the backbone of our approach, guiding the optimisation process for learning the solution distribution of the random differential equation. Section 3 presents the model construction in detail. We define a fundamental representation of Neural Measures tailored to the problem’s probabilistic nature and explore three distinct architectural frameworks, namely fully discrete randomised neural networks, Polynomial Chaos Expansion Neural Networks (PCE-NN), Galerkin-based neural networks. In Section 4, we shift focus to the variational distances used during training, with particular emphasis on the Wasserstein distance and its interpretation in terms of Dirac measures, which naturally relate it to $L^{p}$-type loss functionals. We also discuss alternative metrics and divergences that can be integrated into the framework. Section 5 presents numerical experiments that validate the proposed framework across three representative problems: a bistable ODE, a linear diffusion PDE, and a reaction-diffusion PDE. In each case, the problem parameters are treated as random variables, and the aim is to learn the distribution of the solution. The results demonstrate that the generative PINN models can accurately capture both the mean and higher-order statistical properties of the solutions. Comparisons with reference solvers show very good qualitative and quantitative agreement, though challenges arise in representing sharply peaked distributions and boundary behaviours. Different architectures, including standard PINNs and PINN-PCE variants, are assessed in terms of convergence, error behaviour, and their effectiveness in representing uncertainty. Appendix A presents universal approximation theorems showing that the neural measure spaces, suggested herein, are expressive enough. Therein, we also provide the formal definitions and foundational elements used throughout the work.
Previous works
Physics-Informed Neural Networks (PINNs) are a class of neural network-based methods designed to approximate solutions of partial differential equations (PDEs), where the loss function incorporates the residual of the PDE [41]. Similar approaches have been explored in [24, 5, 40, 45, 31].
To the best of our knowledge, the systematic framework presented here is novel. However, a number of related studies share certain aspects with our approach. A review of neural network approaches to uncertainty quantification can be found in [39]. A method which uses the PINN framework for random differential equations is presented as well. It is assumed the approximate solution is distributed with a normal distribution and the method provides ways to estimate its mean and variance. Inferencing stochastic solution and parameters of a Random PDE, by combining Polynomial Chaos Expansions, [50], and PCA we considered in [58, 57], other related works include [37]. Physics Informed Generative Adversarial Networks were considered in [16] and [53]. See also [32, 21, 48], for UQ related approaches. Neural Operators to address certain issues involving stochasticity in approximating PDEs in infinite dimensional setting include [28, 22, 42, 23, 26, 12, 30], approaches were proposed in [29, 52, 27, 14], their new role and importance in application to sciences was discussed in [15].
2 The problem and its probabilistic framework
2.1 Problem and Assumptions
The objective is to approximate the probability distribution associated with the solution of the following Random (PDE):
(6) $\begin{split}&A_{\xi}u(x,t;\xi)=f(x,t;\xi)\quad\forall\left(x,t\right)\in D\times(0,T]\\ &B_{\xi}u(x,t;\xi)=g(x,t;\xi)\quad\forall\left(x,t\right)\in\Gamma\end{split}$
where $u(x,t;\xi)$ denotes the solution of the PDE, which we interpret as the mapping $U:\Xi\to\mathcal{H}$ with $U(\xi)\coloneqq u(\cdot,\xi)$. Here, $A_{\xi}:\mathcal{H}\to\tilde{\mathcal{H}}$ is a $\xi$-parametrised differential operator, and $B_{\xi}:\mathcal{H}\to\overline{\mathcal{H}}$ is the associated boundary operator. The random parameter $\xi\in\Xi$ is distributed according to a Borel probability measure $\gamma$ on $\Xi$.
Analysing the behaviour of solutions is beyond the scope of this work. Our objective is to propose numerical algorithms for approximating the distribution associated with the solution of the problems at hand. Therefore, we assume that the solution to Eq. 6 exists and is sufficiently regular for our purposes. In particular, we assume throughout that
  • The operators $A_{\xi}$ and $B_{\xi}$ are well-defined for $\gamma$-almost every $\xi$, and the solution of Eq. 6 belongs in $L^{2}(\gamma;\,\mathcal{H})$.
  • For $q\in L^{2}(\gamma;\mathcal{H})$, we assume that the following maps:
    (7) $\begin{split}&\mathcal{A}Q(\xi)\coloneqq A_{\xi}q(\xi),\quad\mathcal{A}Q(\xi)\coloneqq B_{\xi}q(\xi)\qquad\text{ are }L^{2}-\text{integrable, i.e., }\\[3.0pt] &\mathcal{A}Q\in L^{2}(\gamma;\tilde{\mathcal{H}})\quad\text{ and }\quad\mathcal{B}Q\in L^{2}(\gamma;\overline{\mathcal{H}}).\end{split}$
Where the space $L^{2}(\gamma;\mathcal{H})$ is defined in the sense of Bochner integration, see for example Section 7.2.2 of [11] for details. In particular this implies that
$\int\|U(\xi)\|_{\mathcal{H}}^{2}\,\gamma({\rm d}\xi)<\infty.$
and we denote by $U$ the element of $L^{2}(\gamma;\,\mathcal{H})$, corresponding to the solution $u.$ Using the definition of pushforward, we define its probability measure $\nu_{u}\in\mathcal{P}(\mathcal{H})$ by
$\nu_{u}(A)=\left[U\right]_{\!\#}\gamma(A)=\gamma\circ U^{-1}(A),\qquad\forall A\in\mathcal{B}\left(\mathcal{H}\right).$
Since we assume $U\in L^{2}(\gamma;\,\mathcal{H})$, this probability has a finite second moment.
In the next Sections, we introduce a variational method based on a parametrised family of Neural Models to construct an approximation of the target probability distribution $\nu_{u}$.
2.2 Probabilistic framework
We review some standard concepts from probability theory which will be useful in the sequel. We begin by assuming an abstract probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, a measurable space over $\mathcal{H}$, eg. $\left(\mathcal{H},\mathcal{B}\left(\mathcal{H}\right)\right)$, and a random element $\tilde{q}:\Omega\to\mathcal{H}$, i.e., $\mathcal{F}/\mathcal{B}\left(\mathcal{H}\right)$-measurable function. The law of $\tilde{q}$ is the pushforward measure of $\mathbb{P}$ under $\tilde{q}$, defined as:
$\left[\tilde{q}\right]_{\!\#}\mathbb{P}(A)=\mathbb{P}\left(\tilde{q}\in A\right)=\mathbb{P}\circ\tilde{q}^{-1}(A),\qquad A\in\mathcal{B}\left(\mathcal{H}\right),$
where $\tilde{q}^{-1}$ denotes the preimage of the set $A$. It is straightforward to verify that this defines a valid probability measure.
An important assumption in our setting is that we only consider solutions such that $u\in L^{2}\left(\gamma;\,\mathcal{H}\right)$, which ensures two key properties:
  • The mapping $U(\xi)=u(\cdot;\xi)$ is measurable with respect to $\xi$,
  • $U(\xi)$ has a finite second moment with respect to $\gamma$.
Let $\xi:\Omega\to\Xi$ be a random element ( i.e., $\mathcal{F}/\mathcal{B}\left(\Xi\right)$-measurable function) with law $\gamma$. We are interested in random elements $\tilde{q}$ expressible as compositions of a function $q\in L^{2}\left(\gamma;\,\mathcal{H}\right)$ and $\xi$, i.e.,
(8) $\tilde{q}(\omega)=q\left(\xi(\omega)\right),\qquad\mathbb{P}-\text{as }\omega.$
It is also straightforward to verify that $\tilde{q}$ has a finite second moment with respect to $\mathbb{P}$, and its law satisfies:
$\left[\tilde{q}\right]_{\!\#}\mathbb{P}(A)=\left[q\right]_{\!\#}\gamma(A),\qquad\forall A\in\mathcal{B}\left(\mathcal{H}\right).$
Hence setting $\nu_{q}\coloneqq\left[q\right]_{\!\#}\gamma$ we have
$\nu_{q}=\left[q\right]_{\!\#}\gamma\in\mathcal{P}(\mathcal{H}),$
where $\mathcal{P}(\mathcal{H})$ denotes the set of all Borel probability measures on $\mathcal{H}$, i.e., measures defined on $\left(\mathcal{H},\mathcal{B}\left(\mathcal{H}\right)\right)$.
Since our goal is to approximate probability measures, we introduce the following collections:
  • For $p\geq 1$, define the sub-collection $\mathcal{P}_{p}\left(\mathcal{H}\right)$ consisting of all measures with finite $p$-th moment,
    $\mathcal{P}_{p}\left(\mathcal{H}\right)\coloneqq\left\{\begin{split}\mu\in\mathcal{P}\left(\mathcal{H}\right):\,\int_{\mathcal{H}}d_{\mathcal{H}}(h,\overline{h})&{}^{p}\mu({\rm d}h)<\infty\\ &\text{for some (and hence any) }\overline{h}\in\tilde{\mathcal{H}}\end{split}\right\}$,
    where $\left(\mathcal{H},d_{\mathcal{H}}\right)$ is the metric space generating the Borel $\sigma$-algebra $\mathcal{B}\left(\mathcal{H}\right)$.
  • Given a reference measure $\gamma$ on $\mathcal{B}\left(\Xi\right)$, define sub-collection of $\mathcal{P}_{2}\left(\mathcal{H}\right)$
    $\mathcal{P}_{p}^{\gamma}\left(\mathcal{H}\right)=\left\{\mu\in\mathcal{P}(\mathcal{H}):\,\mu=\left[X\right]_{\!\#}\gamma\,\text{ for some }X\in L^{p}\left(\gamma;\mathcal{H}\right)\right\}.$
Over those collections, one may consider alternative metrics for distances or divergences. In this work, we shall use the Wasserstein distance $W_{p}$, [2, 49] and the corresponding $\left(\mathcal{P}_{p}\left(\mathcal{H}\right),W_{p}\right)$ metric space. The $p$-Wasserstein distance is given by
(9) $W_{p}(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\mu,\nu)}\left(\int_{\mathcal{H}\times\mathcal{H}}\,d_{\mathcal{H}}(x,y)^{p}\,\,\pi({\rm d}x,{\rm d}y)\right)^{1/p},$
where $\Pi(\mu,\nu)$ is the set of all plans (joint distributions) with marginals $\mu$ and $\nu$, [2]. Furthermore, we shall use the notation,
(10) $\begin{split}d_{p,\gamma}(\left[X\right]_{\!\#}\gamma,\left[Y\right]_{\!\#}\gamma)^{p}:&=\int_{\mathcal{H}\times\mathcal{H}}d_{\mathcal{H}}(x,y)^{p}\,\left[\left(X,Y\right)\right]_{\!\#}\gamma({\rm d}x,{\rm d}y)\\ &=\int_{\Xi}\left\|X(\xi)-Y(\xi)\right\|_{\mathcal{H}}^{p}\,\gamma({\rm d}\xi).\end{split}$
The distance $d_{p,\gamma}$ induces a metric on $L^{p}(\gamma;\mathcal{H}).$
2.3 Transformed Law due to Parameterised Operators
We now introduce a notation to represent the law of expressions such as $A_{\xi}q(\xi)$ and $B_{\xi}q(\xi)$. In the most general setting, one might define stochastic processes $\tilde{A}:\Xi\times\mathcal{H}\to\tilde{\mathcal{H}}$ and $\tilde{B}:\Xi\times\mathcal{H}\to\overline{\mathcal{H}}$, with appropriate measurability conditions. However, in Section 2.1, we simplify the setting by assuming hypothesis stated in Eq. 7, ensuring measurability. We then define the operation $\odot$ as follows:
$A\odot\left(\left[q\right]_{\!\#}\gamma\right)(C)\coloneqq\left[A_{\xi}q(\xi)\right]_{\!\#}\mathbb{P}(C),$
and similarly for the parameterised operator $B_{\xi}$. For instance we may rewrite the random PDE as equality of probability measures as follows,
[…]
$\mathbb{P}\quad\xrightarrow{\xi}\quad\gamma\quad\begin{cases}\xrightarrow{q}\quad\left[q\right]_{\!\#}\gamma\\ \xrightarrow{A\odot\,}\quad A\odot\,\left[q\right]_{\!\#}\gamma\\ \xrightarrow{B\odot\,}\quad B\odot\,\left[q\right]_{\!\#}\gamma,\end{cases}$
with
$\mathbb{P}\in\mathcal{P}(\Omega),\qquad\gamma\in\mathcal{P}(\Xi),\quad\left[q\right]_{\!\#}\gamma\in\mathcal{P}_{2}^{\gamma}(\mathcal{H})$
$A\odot\,\left[q\right]_{\!\#}\gamma\in\mathcal{P}_{2}^{\gamma}(\tilde{\mathcal{H}})\quad\text{and}\quad A\odot\,\left[q\right]_{\!\#}\gamma\in\mathcal{P}_{2}^{\gamma}(\overline{\mathcal{H}}).$
2.4 Variational Approximation of the Target Law
To approximate the target probability measure $\nu_{u}$, we adopt a variational approach. Specifically, we minimise a discrepancy functional over $\mathcal{V}_{\Theta}=\mathcal{V}_{\mathcal{N}}^{*}$ that quantifies the difference between transformed distributions of the model and the target, induced by the operators $A_{\xi}$ and $B_{\xi}$. Alternative specific choices for the neural measure spaces $\mathcal{V}_{\Theta}=\mathcal{V}_{\mathcal{N}}^{*}$ are provided in the next section. The discrepancy we minimise is given by:
$d_{\mathcal{P}_{2}^{\gamma}(\tilde{\mathcal{H}})}\left(A\odot\nu_{u},A\odot\mu_{\theta}\right)+d_{\mathcal{P}_{2}^{\gamma}(\overline{\mathcal{H}})}\left(B\odot\nu_{u},B\odot\mu_{\theta}\right),$
[…]
(11) $\mu_{\theta^{*}}\in\mathop{\mathrm{arg\,min}}_{\mu_{\theta}\in\mathcal{V}_{\mathcal{N}}^{*}}d_{\mathcal{P}_{2}^{\gamma}(\tilde{\mathcal{H}})}(A\odot\nu_{u},A\odot\mu_{\theta})+d_{\mathcal{P}_{2}^{\gamma}(\overline{\mathcal{H}})}(B\odot\nu_{u},B\odot\mu_{\theta}).$
By varying the choice of metrics, neural architectures $\mathcal{V}_{\mathcal{N}}$ and the structure of $X_{\theta}$, different approximation strategies can be implemented. Nonetheless, the ultimate objective remains the same: to identify an optimal Neural Measure $\mu_{\theta^{*}}$ that is close enough to the target distribution $\nu_{u}$.
3 Discrete Neural Measure Spaces
To motivate the definition of neural measures on $\mathcal{P}(\mathcal{H})$ we consider a family of discrete stochastic mappings
$\tilde{X}_{\theta}:\Omega\to\mathcal{H},$
where each map $\tilde{X}_{\theta}$ is associated with a neural network function parameterised by $\theta\in\Theta$, where $\Theta$ indexes a class of neural networks $\mathcal{V}_{\mathcal{N}}$. Our building assumption is that $\tilde{X}_{\theta}$ will be based on standard discrete neural spaces $\mathcal{V}_{\mathcal{N}}$ designed to approximate functions. The alternative designs we consider are based on neural network functions corresponding to (i) both deterministic ($x,t$) and stochastic variables ($\xi$), (ii) only deterministic variables and (iii) only stochastic variables. Thus, each model will be characterised by a function $g_{\theta}\in\mathcal{V}_{\mathcal{N}}$, and our goal is to approximate the distribution associated to the solution $u$ by solving an appropriate optimisation problem over $\Theta$.
To be more precise, let $\mathcal{V}_{\mathcal{N}}$ be a chosen class of neural networks as above, and let $\gamma$ be the reference probability measure defined on a latent space $\Xi$. We define first a corresponding family of functions $X_{\theta}=X_{\theta}(x,t,\xi),$
$X_{\theta}:\Xi\to\mathcal{H},$
where each model $X_{\theta}$ is implicitly determined by a function $g_{\theta}\in\mathcal{V}_{\mathcal{N}},$ i.e., $X_{\theta}=X(g_{\theta}).$ These functions encode the neural architecture and the parameterisation of the model, and will be detailed in the following Sections 3.1-3.
To introduce stochasticity, we consider a random variable $\xi:\Omega\to\Xi$, distributed according to $\gamma$. Composing this with the deterministic map $X_{\theta}$ yields the stochastic model:
$\tilde{X}_{\theta}(\omega)=X_{\theta}\circ\xi(\omega).$
This composition defines a random element in $\mathcal{H}$ allowing us to generate samples from the associated distribution. The pushforward of $\gamma$ through $X_{\theta}$, denoted by $\left[X_{\theta}\right]_{\!\#}\gamma$, defines the induced discrete Neural Measure Space on $\mathcal{P}(\mathcal{H}).$ In fact, the class of Neural Measures spaces we consider have the form,
$\mathcal{V}_{\mathcal{N}}^{*}=\left\{\mu_{\theta}=\left[X_{\theta}\right]_{\!\#}\gamma\text{ for some Neural Model }X_{\theta}=X(g_{\theta}),\,\ \ g_{\theta}\in\mathcal{V}_{\mathcal{N}}\,\right\}.$
Clearly, $\mathcal{V}_{\mathcal{N}}^{*}\subset\mathcal{P}(\mathcal{H})$ and in fact $\mathcal{V}_{\mathcal{N}}^{*}\subset\mathcal{P}_{2}^{\gamma}\left(\mathcal{H}\,\right).$ Here, $\mathcal{V}_{\mathcal{N}}$ is fixed a priori, and the notation $\mathcal{V}_{\mathcal{N}}^{*}$ reflects the fact that this space is associated with a discrete standard neural network space $\mathcal{V}_{\mathcal{N}}.$ Each model in $\mathcal{V}_{\mathcal{N}}^{*}$ corresponds to a unique pushforward measure derived from a neural map and the reference measure. Notice finally that although $\xi\sim\gamma$ is sampled from a known reference distribution, its explicit form $\xi\left(\omega\right)$ need not to be specified.
3.1 Fully Network Based Neural Measure Spaces
Next, we consider $X_{\theta}$, defined in relation to neural network functions $g_{\theta}\in\mathcal{V}_{\mathcal{N}}$, corresponding to both deterministic ($x,t$) and stochastic variables ($\xi$). The resulting neural models solely rely on a chosen class of neural networks $\mathcal{V}_{\mathcal{N}}$ without any further approximation step.
By fixing the number of layers, the size of each hidden layer, and the activation functions, we define the following class
$\mathcal{V}_{\mathcal{N}}\coloneqq\left\{\begin{aligned} g_{\theta}:\mathbb{R}^{d_{\text{in}}}\times\mathbb{R}^{d_{\text{ran}}}&\to\mathbb{R}^{s}\\ (x,t;\xi)&\mapsto(C_{L}\circ\sigma_{L-1}\circ\dots\circ\sigma_{1}\circ C_{1})(x,t;\xi)\end{aligned}\;\;:\;\;\theta\in\Theta\right\}.$
In the above procedure, any such map $g_{\theta}$ is defined by the intermediate layers $C_{k}$, which are affine maps of the form
(12) $C_{k}y=W_{k}y+b_{k},\quad\text{with}\quad W_{k}\in\mathbb{R}^{d_{k+1}\times d_{k}},\;\;b_{k}\in\mathbb{R}^{d_{k+1}},$
where the dimensions $d_{k}$ may vary with each layer $k$ and the activation function $\sigma(y)$ denotes the vector with the same number of components as $y$, where $\sigma(y)_{i}=\sigma(y_{i}).$ The index $\theta\in\Theta$ represents collectively all the parameters of the network.
Then, we define the neural model as:
$X_{\theta}(\xi)=g_{\theta}\left(\cdot;\xi\right),\text{ for some }g_{\theta}\in\mathcal{V}_{\mathcal{N}}.$
We may additionally require that for every $g_{\theta}\in\mathcal{V}_{\mathcal{N}}$
  • $g_{\theta}(\cdot,\xi)\in\mathcal{H}$, for every $\xi$,
  • and $\int_{\Xi}\left\|g_{\theta}(\cdot;\xi)\right\|_{\mathcal{H}}^{2}\gamma({\rm d}\xi)<\infty$.
These ensure that $X_{\theta}\in L^{2}(\gamma;\mathcal{H})$. The associated variational method is optimised based on the following class of Neural Measures:
$\mathcal{V}_{\mathcal{N}}^{*}=\left\{\left[X_{\theta}\right]_{\!\#}\gamma\in\mathcal{P}_{p}^{\gamma}\left(\,\,\mathcal{H}\,\,\right):\text{ where }X_{\theta}(\xi)=g_{\theta}(\cdot,\xi)\text{ for some }g_{\theta}\in\mathcal{V}_{\mathcal{N}}\right\}.$
This formulation is simple and general–it captures a wide range of model classes without assuming any structure other than that provided by the neural network architecture itself. The construction presented is typical, and several alternative architectures are feasible; it is presented in this specific form for completeness of exposition.
3.2 PCE-NN Based Neural Measure Spaces
In this section, we introduce a class of neural models designed to approximate functions $U\in L^{2}(\gamma;\mathcal{H})$, where the computational complexity of learning high-dimensional mappings $\left(x,t;\xi\right)\mapsto\mathbb{R}^{s}$ is reduced through truncated spectral expansion. The key idea is to exploit the linear structure of the Hilbert space $L^{2}(\gamma;\mathcal{H})$ and separate the stochastic and deterministic components of the function.
To this end, consider a Polynomial Chaos Expansion (PCE)-type basis, [50], $\left\{\phi_{n}\right\}_{n\in\mathbb{N}}\subset L^{2}(\gamma)$ satisfying:
(13) $\forall q\in L^{2}(\gamma;\mathcal{H}),\quad\exists\,!\,\left\{a_{n}\right\}_{n\in\mathbb{N}}\subset\mathcal{H}:\quad q(x,t;\xi)=\sum_{n\in\mathbb{N}}a_{n}(x,t)\phi_{n}(\xi)\vskip-8.5359pt,$
we represent the neural model as a truncated expansion:
$X_{\theta}(\xi)=\sum_{n=1}^{K}g_{\theta}^{(n)}\,\phi_{n}(\xi),$
where each coefficient function $g_{\theta}^{(n)}\in\mathcal{H}$ is a component of the output layer of some neural network $g_{\theta}$ within $\mathcal{V}_{\mathcal{N}}$, i.e.
(14) $\mathcal{V}_{\mathcal{N}}\coloneqq\left\{g_{\theta}:\mathbb{R}^{d_{\text{in}}}\to\mathbb{R}^{d_{\text{out}}K}:\ g_{\theta}=g_{\theta}(x,t)\text{ for }\theta\in\Theta\right\},$
and the structure of the neural network functions $g_{\theta}(x,t)$ is as before, but now affecting only $x,t$ variables. Then, the corresponding form of neural measures is:
(15) $\mathcal{V}_{\mathcal{N}}^{*}=\left\{\left[X_{\theta}\right]_{\!\#}\gamma\in\mathcal{P}_{2}^{\gamma}\left(\,\,\mathcal{H}\,\,\right):\text{ where }X_{\theta}(\xi)=\sum_{n=1}^{K}g_{\theta}^{(n)}\,\phi_{n}(\xi)\text{ for some }g_{\theta}\in\mathcal{V}_{\mathcal{N}}\right\}.$
This construction decouples the stochastic and deterministic components, enabling: simpler and lower-dimensional neural architectures, and possibly structured learning problems.
3.3 Galerkin-NN Based Neural Measure Spaces
Galerkin-NNs are conceptually similar to PCE-PINNs but draw inspiration from the Galerkin method used in numerical PDEs. In this approach, the governing equations are projected onto a test basis, and neural networks are formulated accordingly. The architecture closely mirrors that of PCE-NNs, although neural networks are used now for the discretisation of $\xi$ variable. For this reason, we only present the analogue corresponding to the variant that directly approximates the solution $U(\xi)$.
Consider appropriate a given finite dimensional Galerkin subspace of $\mathcal{H}$, denoted by $\mathcal{H}_{h},$ let $M$ be its dimension and $\left\{\psi_{n}\right\}_{1}^{M}$ its basis. Then, we consider the elements of $L^{2}(\gamma;\mathcal{H})$
(16) $q(x,t;\xi)=\sum_{n=1}^{M}b_{n}(\xi)\psi_{n}(x,t).$
Clearly, for each fixed $\xi,$ $q(\cdot;\xi)\in\mathcal{H}_{h}.$ We then represent the neural model as
$X_{\theta}(\xi)=\sum_{n=1}^{M}g_{\theta}^{(n)}\left(\xi\right)\,\psi_{n},$
where each coefficient function $g_{\theta}^{(n)}\in L^{2}(\gamma)$ is a component of the output layer of some neural network $g_{\theta}$ within $\mathcal{V}_{\mathcal{N}}$, i.e.,
$\mathcal{V}_{\mathcal{N}}\coloneqq\left\{g_{\theta}:\mathbb{R}^{d_{\text{ran}}}\to\mathbb{R}^{M}:g_{\theta}=g_{\theta}(\xi)\text{ for }\theta\in\Theta\right\},$
and the structure of the neural network functions $g_{\theta}(\xi)$ is as before, but now affecting only the $\xi$ variable. Then, the corresponding form of neural measures is:
$\mathcal{V}_{\mathcal{N}}^{*}=\left\{\left[X_{\theta}\right]_{\!\#}\gamma\in\mathcal{P}_{2}^{\gamma}\left(\,\,\mathcal{H}\,\,\right):\text{ where }X_{\theta}(\xi)=\sum_{n=1}^{M}g_{\theta}^{(n)}\left(\xi\right)\,\psi_{n}\text{ for some }g_{\theta}\in\mathcal{V}_{\mathcal{N}}\right\}.$
Although the above formulation is quite flexible and compatible with more standard methods for the numerical solution of PDEs, its actual implementation depends on the choice of the spaces $\mathcal{H}_{h}.$ It is possible using the Discontinuous Galerkin framework, to consider spaces that are not subspaces of $\mathcal{H}$. This approach is more involved and is beyond the scope of this work.
4 Distances and their evaluation
In statistical learning and probabilistic modeling, measuring the ”distance” between probability measures is a fundamental task. Two broad families of such distances are commonly used: divergences and metrics, both of which serve to quantify dissimilarity over a suitable space of probability measures, such as $\mathcal{P}(\mathcal{H})$. Depending on the nature of the problem, several alternative choices can be made, resulting in different loss functions. Our framework is flexible in this regard. We will focus next on Wasserstein distances, as in this case, the loss function takes a particular simple form.
4.1 Loss based on Wasserstein distance
Within the framework the present paper it is particularly convenient to consider Wasserstein distances. A key observation is that when one of the measures to be used in the loss is atomic, i.e., $\delta_{g},$ then the set of plans is reduced to $\mu\times\delta_{g},$ i.e., $\Pi(\mu,\delta_{g})=\{\mu\times\delta_{g}\}$, see (5.2.12) in [2]. Assuming $g\in\mathcal{H}$ and $\delta_{g}\in\mathcal{P}_{p}(\mathcal{H})$, we then have
(17) $\begin{split}W_{p}(\mu,\delta_{g})^{p}=&\inf_{\pi\in\Pi(\mu,\delta_{g})}\int_{\mathcal{H}}\int_{\mathcal{H}}\left\|x-y\right\|_{\mathcal{H}}^{p}\pi({\rm d}x,{\rm d}y)\\ =&\int_{\mathcal{H}}\int_{\mathcal{H}}\left\|x-y\right\|_{\mathcal{H}}^{p}\mu({\rm d}x)\delta_{g}({\rm d}y)\\ =&\int_{\mathcal{H}}\left\|x-g\right\|_{\mathcal{H}}^{p}\mu({\rm d}x)=\mathbb{E}_{X\sim\mu}\left\|X-g\right\|_{\mathcal{H}}^{p}.\end{split}$
We distinguish two cases. In case where $\mathcal{A}U(\xi)$ and $\mathcal{B}U(\xi)$ are constant with respect to $\xi$, we can easily check exists $f\in\tilde{\mathcal{H}}$ and $g\in\overline{\mathcal{H}}$ such that
$A\odot\nu_{u}=\delta_{f},\qquad B\odot\nu_{u}=\delta_{g},$
and the suggested variational method can be written as follows:
$\mu_{\theta^{*}}\in\operatorname*{arg\,min}_{\mu_{\theta}\in\mathcal{V}_{\mathcal{N}}^{*}}W_{p}(A\odot\mu_{\theta},\delta_{f})+W_{p}(B\odot\mu_{\theta},\delta_{g})$
In the general case, it is still possible to use this property if we define residual operators
$\tilde{A}_{\xi}X_{\theta}(\xi)\coloneqq\mathcal{A}X_{\theta}(\xi)-\mathcal{A}U(\xi),\quad\text{ and }\quad\tilde{B}_{\xi}X_{\theta}(\xi)\coloneqq\mathcal{B}X_{\theta}(\xi)-\mathcal{B}U(\xi),$
and the suggested variational method can be written as follows:
$\mu_{\theta^{*}}\in\operatorname*{arg\,min}_{\mu_{\theta}\in\mathcal{V}_{\mathcal{N}}^{*}}\,\,W_{p}(\tilde{A}\odot\mu_{\theta},\delta_{0})+W_{p}(\tilde{B}\odot\mu_{\theta},\delta_{0}).$
This is particularly helpful, since the computation of the Wasserstein distances in both cases is reduced to ($p=2$) the evaluation of integrals of functions,
$\begin{split}d_{2,\gamma}\left(A\odot\nu_{u},\,A\odot\mu_{\theta}\right)^{2}+&d_{2,\gamma}\left(B\odot\nu_{u},\,B\odot\mu_{\theta}\right)^{2}\\ &=\left\|\mathcal{A}U-\mathcal{A}X_{\theta}\right\|_{L^{2}(\gamma;\tilde{\mathcal{H}})}^{2}+\left\|\mathcal{B}U-\mathcal{B}X_{\theta}\right\|_{L^{2}(\gamma;\overline{\mathcal{H}})}^{2}.\end{split}$
[…]
4.2 Other metrics and divergences
It is possible, however, to utilise alternative divergences or metrics to define the loss functional. We briefly review some potential choices below, see, eg., [7, 43, 33, 46, 6, 49, 44].
4.2.1 Divergences and Metrics: Definitions and Relationships
Divergences generalise metrics by dropping symmetry and the triangle inequality, while retaining non-negativity and vanishing only when the inputs are equal. A key family is the f-divergences: given probability measures $\mu\ll\nu$ and a convex function $f:\mathbb{R}_{+}\to\mathbb{R}\cup\left\{+\infty\right\}$ with $f(1)=0$, the $f$-divergence is defined as
$D_{f}(\mu||\nu)=\int f\left(\frac{{\rm d}\mu}{{\rm d}\nu}\right){\rm d}\nu.$
This includes the Kullback–Leibler (KL) divergence, Jensen–Shannon divergence, and others. In addition to divergences, several frameworks exist for constructing metric distances between probability measures. These include:
  • (A)
    Integral Probability Metric (IPM):
    Given a class of test functions $\mathcal{F}$, the IPM between two measures $\mu,\,\nu$ is defined as:
    (18) $d_{\mathcal{F}}(\mu,\nu)\coloneqq\sup_{f\in\mathcal{F}}\left|\mathbb{E}_{X\sim\mu}f(X)-\mathbb{E}_{Y\sim\nu}f(Y)\right|.$
  • (B)
    Optimal Transport (OT) Metrics:
    In OT theory, distances are defined via couplings. Given a cost function $c(x,y)$, the Wasserstein distance (or more generally, $c$-Wasserstein distance) is
    $W_{c}(\mu,\nu)=\inf_{\pi\in\Pi(\mu,\nu)}\int c(x,y){\rm d}\pi(x,y),$
    where $\Pi(\mu,\nu)$ is the set of couplings of $\mu$ and $\nu$.
  • (C)
    TV Distance via Densities:
    Given a probability measure $\lambda\in\mathcal{P}(\mathcal{H})$, we define $\mathcal{P}_{\lambda}(\mathcal{H})\subset\mathcal{P}(\mathcal{H})$ to be the collection of probability measures that are absolutely continuous with respect to $\lambda$, i.e., $\mu\ll\lambda$. Let $\mu,\,\nu\in\mathcal{P}_{\lambda}(\mathcal{H})$ with Radon-Nikodym derivatives $f_{\mu},\,f_{\nu}$ with respect to $\lambda$. Then, TV distance between $\mu,\,\nu$ can be expressed as:
    $d_{TV}(\mu,\nu)\coloneqq\frac{1}{2}\left\|f_{\mu}-f_{\nu}\right\|_{L^{1}(\lambda)}.$
4.2.2 Approximating Distances in Finite-Dimensional Spaces
Given that our spaces, $V_{\Theta}=V_{\mathcal{N}}^{*}$, have a finite number of parameters, we can employ techniques developed in approximating distances of distributions in finite dimensions to compute the actual loss function. Indicative such methods include, GANs and computational optimal transport algorithms, [17, 36].
For approximating Wasserstein distances, one may use Sinkhorn Algorithm which introduces entropic regularisation for efficient computation and other approaches detailed in [36]. These approaches exploit the geometric and variational structure of optimal transport theory and are particularly relevant in generative modelling and distributional approximation tasks.
Generative Adversarial Networks (GANs) provide a flexible variational framework for approximating probability distributions by minimising statistical distances, notably Integral Probability Metrics (IPMs) and, under certain conditions, $f$-divergences.
Given a class of generators $g_{\theta}\in\mathcal{V}_{\mathcal{N}}$, a class of discriminators $f_{\phi}\in\overline{\mathcal{V}}_{\mathcal{N}}$, a reference distribution $\pi_{0}$ and target distribution $\mathbb{P}_{r}$ the standard GAN objective takes the form:
$\min_{g_{\theta}\in\mathcal{V}_{\mathcal{N}}}\max_{f_{\phi}\in\overline{\mathcal{V}}_{\mathcal{N}}}\,\,\mathbb{E}_{X\sim\mathbb{P}_{r}}f_{\phi}(X)-\mathbb{E}_{Z\sim\pi_{0}}f_{\phi}\circ g_{\theta}\circ(Z).$
In the special case where the class of test functions $\mathcal{F}$ associated with a particular IPM (eg. as in equation Eq. 18) can be approximated by the class of discriminators $\overline{\mathcal{V}}_{\mathcal{N}}$. Then, this min-max formulation can be viewed as minimising IPM between $\mathbb{P}_{r}$ and generators distribution. Specifically, we define the class of neural measures:
[…]
where $\left[g_{\theta}\right]_{\!\#}\pi_{0}$ is the law of the distribution $g_{\theta}\circ(Z)$ when $Z\sim\pi_{0}$. The above min-max formulation provides a way to approximate the variational problem:
[…]
5 Numerical results
In this section, we present numerical experiments that illustrate the performance of the methods introduced in Sections 2 and 3. We consider three representative examples: a bistable ordinary differential equation (ODE), a diffusion partial differential equation (PDE), and a reaction-diffusion PDE. In all cases, the problem parameters are treated as random variables drawn from prescribed distributions, and the objective is to quantify the resulting uncertainty in the solution.
All neural networks employed are fully connected multi-layer perceptrons (MLPs). The architectures used in the three examples consist of 5, 6, and 6 hidden layers, with 32, 20, and 40 hidden units per layer, respectively. We use the sinusoidal-based activation function known as the snake activation, defined as
$\sigma(x;a)=x+\frac{\sin^{2}(ax)}{a},$
following [60]. This choice consistently yielded better convergence and expressiveness than the traditional hyperbolic tangent activation, especially for problems involving both periodic and non-periodic solutions. The parameter $a$ is specific to each layer and is optimized jointly with the network weights during training.
Training is performed using the L-BFGS optimizer with a learning rate of 1, a maximum number of iterations set to 20, and a history size of 20. No batch processing is used; the optimizer is applied to the full dataset at each step. We initialize the weights using Xavier initialization. Training samples from the spatiotemporal domain of the differential equation are resampled every predefined number of iterations, while samples from the parameter space are resampled every different predefined interval. The resampling frequency for the domain is higher than that for the parameter space. This strategy allows the network to first focus on learning the PDE structure for a specific parameter configuration and later generalize to new configurations as training progresses.
All experiments were conducted using a unified codebase, which is publicly available at [3], ensuring reproducibility and ease of extension to other problem settings.
5.1 Bistable ordinary differential equation
We consider the ordinary differential equation,
(19) $\frac{\mathrm{d}u}{\mathrm{d}t}=-r(u-1)(2-u)(u-3),\quad t\in[0,8],$
with initial condition,
(20) $u(0)=u_{0}.$
The parameters $u_{0},r$ follow uniform distributions, $u_{0}\sim\mathcal{U}(0,4)$ and $r\sim\mathcal{U}(0.8,1.2)$. The equation has two stable ($u=1$, $u=3$) and one unstable ($u=2$) equilibrium solutions.
In Figure 4, we plot the training and testing loss functions throughout the optimisation. The spatial-temporal domain is resampled every 50 iterations, as are the parameters, which are drawn from their predefined uniform distribution. Each resampling uses 100 random samples from the domain. After each resampling step, the loss typically spikes due to the shift in training data but continues to decrease overall as training progresses. The testing loss begins to rise after around 400 iterations, indicating that the network starts to overfit the training data.
In Figure 1, we plot histograms of the solution to the bistable ODE Eq. 19. The reference solution is obtained using an implicit numerical solver. We observe good qualitative agreement between the reference and the PINN-based histograms. However, the Wasserstein distance between the two solutions increases over time. This is expected, as the solution’s distribution gradually concentrates into two Dirac delta peaks at $u=1$ and $u=3$. This sharp concentration makes convergence challenging, since the Wasserstein distance is highly sensitive to samples that have not fully collapsed onto the Dirac masses.
Figure 1: Histogram comparison at time $t=2$ of the solution of the bistable ODE Eq. 19 using an implicit numerical solver and the PINN architecture for different parameter values.
5.2 Diffusion equation
We solve the diffusion equation
(21) $u_{t}-\frac{a}{k^{2}}u_{xx}=0,\quad t\in[0,1],\quad x\in[0,\pi],$
with initial and boundary conditions,
(22) $u(0,x)$ $=\sin(kx),\quad x\in[0,\pi],$
(23) $u(t,0)$ $=0,\quad t\in[0,1],$
(24) $u(t,\pi)$ $=e^{-at}\sin(\pi k),\quad t\in[0,1],$
and exact solution,
(25) $u(t,x)=\mathrm{e}^{-kt}\sin(\pi x).$
We assume that the parameters $a,k$ follow uniform distributions, $a\sim\mathcal{U}(1,3)$ and $k\sim\mathcal{U}(1,3)$.
In Figure 5, we plot the training and testing loss functions during the training of both the PINN and PINN-PCE architectures for the diffusion equation Eq. 21. The spatial-temporal domain is resampled every 50 iterations, while the parameters are resampled every 100 iterations. The domain is sampled by selecting 20 spatial and 10 temporal points, forming a Cartesian product of 200 training locations. The PCE basis used in the PINN-PCE model consists of orthogonal polynomials up to degree 5. We observe that the training and testing errors of the PINN-PCE model are consistently higher than those of the plain PINN model. This behavior remains even when increasing the degree of the PCE basis.
(a) Mean and standard deviation of the reference and PINN-based solutions.
(b) Average values comparison.
(c) Histogram comparison at time $t=0.5$.
Figure 2: Average and one standard deviation of the reference and PINN-based solutions (a), at time $t=0.5$ and $t=1$. Average values (b) and histograms of the solution (c) of the diffusion equation Eq. 21 at time $t=0.5$ using the PINN and the PINN-PCE architectures. Notice that PINN produces marginally better results than the PINN-PCE architecture.
In Figure 2, we present the average solution and histograms for the diffusion equation Eq. 21 at time $t=0.5$, computed using the PINN and PINN-PCE architectures. In Figure 2a, the x-axis represents the spatial coordinate and the y-axis corresponds to time. Figure 2b displays the spatial coordinate on the x-axis and the logarithm of the frequency of solution values on the y-axis. We observe that the PINN-PCE architecture yields slightly improved results compared to the plain PINN. Both methods show good qualitative agreement with the analytical solution, indicating their capacity to learn the underlying dynamics of the equation.
5.3 Reaction-Diffusion
We consider the partial differential equation,
(26) $u_{t}-Du_{xx}+gu^{3}=f,\quad t\in[0,4],\ x\in[-1,1],$
with initial and boundary conditions,
(27) $u(0,x)=0.5\cos^{2}(\pi x),$
(28) $u(t,-1)=u(t,1)=0.5.$
The reaction function $g$ is given by,
(29) $g(x)=0.2+e^{r_{1}x}\cos^{2}(r_{2}x),$
with $r_{1}\sim\mathcal{U}(0.5,1)$ and $r_{2}\sim\mathcal{U}(3,4)$ and the forcing function is given by,
(30) $f(x)=\exp\left(-\frac{(x-0.25)^{2}}{2k_{1}^{2}}\right)\sin^{2}(k_{2}x),$
with $k_{1}\sim\mathcal{U}(0.2,0.8)$ and $k_{2}\sim\mathcal{U}(1,4)$. The diffusion coefficient is fixed to $D=0.01$. This equation has been studied in [39].
In Figure 6, we plot the training and testing loss functions during the training of the PINN architecture for the reaction-diffusion equation Eq. 27. The spatial-temporal domain is resampled every 50 iterations, while the parameters are resampled every 100 iterations. The domain is sampled by selecting 40 spatial and 20 temporal points, forming a Cartesian product of 800 training locations.
(a) Mean and standard deviation of the reference and PINN-based solutions.
(b) Absolute error heatmap.
(c) Histograms of the solution at time $t=2$.
Figure 3: Average values and histograms of the solution of the reaction-diffusion equation Eq. 27 at time $t=2$ using the PINN architecture. The statistics produced by the PINN architecture are in good qualitative agreement with the reference solution.
In Figure 3, we plot the mean and standard deviation of the reference and PINN-based solutions. The absolute error heatmap shows that the PINN-based solution is in good qualitative agreement with the reference solution. The histograms of the solution at $t=2$ show that the PINN-based solution is in good qualitative agreement with the reference solution. The error shown in Figure 3b shows larger values at the boundaries of the domain for initial times. This flaw can probably be fixed by using adaptive selection of the training points based on the error.
Appendix A Universal Approximation Theorem
Universal Approximation Theorems (UATs) are foundational results in the theory of neural networks. They establish that certain classes of functions—often defined by specific neural network architectures—are dense in a target function space. In other words, these theorems guarantee that the model class is expressive enough to approximate any function in the target space to an arbitrary degree of accuracy.
The classical Universal Approximation Theorem states that neural networks with a continuous, non-polynomial activation function are dense in the space of continuous functions on compact subsets $\mathbb{R}^{d}$ [10, 18, 38]. Subsequent developments extended these results to approximation in Sobolev spaces and provided quantitative approximation rates for specific activations such as $\text{ReLU},$ $\text{ReLU}^{k}$ and $\tanh$ [54, 35, 13, 51]. However, in many applications – and in particular for parametric PDEs – the relevant function domains are non-compact, and the target functions take values in infinite-dimensional Banach spaces such as Sobolev spaces. In this direction, Neufeld and Schmocker [34] proved a universal approximation theorem on non-compact domains for networks with general non-polynomial activations, providing density results in appropriate weighted function spaces. Meanwhile, universal approximation results for operator-learning networks mapping between Banach-valued function spaces have been established in more specialised settings, such as Theorem 9.3 in Kovachki et al. [22], which applies to certain classes of neural operators.
Building on these developments, we show below that the neural-network-based spaces of probability measures we consider are capable of approximating probability measures represented through Bochner spaces $L^{2}(\gamma;\mathcal{H})$, where $\mathcal{H}$ is a general Hilbert space (eg., $H^{s}(D)$). Thus, our Neural Measures spaces are expressive enough to yield accurate approximations for our purpose.
Revisiting UAT for Banach Spaces $X$
Let $U$ a Borel subset of $\mathbb{R}^{m}$, and denote by $C_{b}^{k}(U;\mathbb{R}^{d})$ the vector space of $k$-times continuously differentiable functions $f:U\to\mathbb{R}^{d}$ such that both the function and all its partial derivatives up to order $k$ are continuous and bounded. In the case $k\geq 1$, we additionally assume that $U$ is open.
For a multi-index $a=\left(a_{1},\dots a_{m}\right)$ with $|a|\leq k$, the $a$-th partial derivative is given by
(31) $u\mapsto\partial_{a}f(u)\coloneqq\frac{\partial^{|a|}}{\partial u_{1}^{a_{1}}\dots\partial u_{m}^{a_{m}}}f(u)\in\mathbb{R}^{d}.$
Equipped with the norm
$\left\|f\right\|_{C_{b}^{k}(U;\mathbb{R}^{d})}\coloneqq\max_{|a|\leq k}\,\sup_{u\in U}\left|\partial_{a}f(u)\right|,$
the space $C_{b}^{k}(U;\mathbb{R}^{d})$ becomes a Banach space.
To introduce weighted variants, we define the polynomially weighted $C^{k}$-norm:
$\left\|f\right\|_{C_{pol,\beta}^{k}(U;\mathbb{R}^{d})}\coloneqq\max_{|a|\leq k}\,\sup_{u\in U}\frac{\left|\partial_{a}f(u)\right|}{\left(1+\left|u\right|\right)^{\beta}},\qquad\text{ for some }\beta\geq 0.$
Using this norm, we define the weighted $C^{k}$-space $\overline{C_{b}^{k}(U;\mathbb{R}^{d})}^{\beta}$ as the closure of $C_{b}^{k}(U;\mathbb{R}^{d})$ under the $\left\|\cdot\right\|_{C_{pol,\beta}^{k}(U;\mathbb{R}^{d})}$-norm. In either case (bounded, or unbounded $U$), this space is naturally identified as subset of $C_{b}^{k}(U;\mathbb{R}^{d})$; for details, see Lemma 4.1 of [34]. Thus, equipped with the weighted norm, $\left(\overline{C_{b}^{k}(U;\mathbb{R}^{d})}^{\beta},\left\|\cdot\right\|_{C_{pol,\beta}^{k}(U;\mathbb{R}^{d})}\right)$, is Banach space. {assumption}[for UAT in Banach Spaces] Let $k\geq 0,\,U\subset\mathbb{R}^{m},$ and $\beta>0$. In the special case where $k\geq 1$, we assume $U$ is open. Let $\left(X,\left\|\cdot\right\|_{X}\right)$ be a Banach space of functions $f:U\to\mathbb{R}^{m}$ such that the restriction map
$\left(C_{b}^{k}(\mathbb{R}^{m};\mathbb{R}^{d}),\left\|\cdot\right\|_{C_{pol,\beta}^{k}}\right)\ni f\mapsto f\Big|_{U}\in\left(X,\left\|\cdot\right\|_{X}\right),$
is a continuous dense embedding111 “Continuous dense embedding” means the restriction map $i_{U}\!\!:\!\!\left(\!C_{b}^{k}(\!\mathbb{R}^{m};\!\mathbb{R}^{d}\!),\!\left\|\cdot\right\|_{C_{pol,\beta}^{k}}\!\!\right)\!\!\to\!\!\left(\!X,\!\left\|\cdot\right\|_{X}\!\right)$ is a bounded linear injection with dense image, i.e., $\|i_{U}\|<\infty$ and $\operatorname{cl}_{\left\|\cdot\right\|_{X}}\left(Im(i_{U})\right)=X$. .
Theorem A.1 (Restating Theorem 2.7 in [34], UAT for approximating Banach space $X$).
Suppose we are given parameters $k,\,\beta$, a Borel subset $U\subset\mathbb{R}^{m}$ and a Banach space $X$ consisting of functions with domain $U$ such that Appendix A holds. Let $\rho\in\overline{C_{b}^{k}(\mathbb{R})}^{\beta}$ be a non-polynomial activation function, and consider the following class of neural networks:
(32) $\mathcal{N}_{\rho,\mathbb{R}^{m},d}^{(K)}=\left\{\begin{split}\phi:\mathbb{R}^{m}\to\mathbb{R}^{d}\text{ with }\phi(u)=\sum_{n=1}^{K}y_{n}\rho(a_{n}^{T}u-b_{n}),&\\ &~\hskip-170.71652pt\text{ for some }a_{1},\dots a_{K}\in\mathbb{R}^{m},\,b_{1},\dots b_{K}\in\mathbb{R},\,y_{1},\dots y_{K}\in\mathbb{R}^{d}\end{split}\right\},$
and let $\mathcal{N}_{\rho,\mathbb{R}^{m},d}\coloneqq\bigcup_{K\in\mathbb{N}}\mathcal{N}_{\rho,\mathbb{R}^{m},d}^{(K)}$. Then, it holds that
$cl_{\left\|\cdot\right\|_{X}}\left(\mathcal{N}_{\rho,\mathbb{R}^{m},d}\Big|_{U}\right)=X,$
i.e., the closure of the neural network class $\mathcal{N}_{\rho,\mathbb{R}^{m},d}$ (restricted to $U$) with respect to the $X$-norm coincides with $X$ itself.
Several examples of Banach spaces satisfying Appendix A are provided in Example 2.6 of [34].
UAT for PCE-NN/ Galerkin-NN Based Neural Measure Spaces
In both versions of the Universal Approximation Theorem (UAT) for Neural Measure Space architectures introduced in Sections 3.2 and 3.3, we work within the Bochner space $L^{2}(\gamma;\mathcal{H})$. In each case, the approximation results rely on expansions of this space: either in a basis of $L^{2}(\gamma)$ or in a basis of $\mathcal{H}$. More generally, such expansions may also be formulated directly in a basis of $L^{2}(\gamma;\mathcal{H})$ itself.
A sufficient condition ensuring the existence of these bases is the separability of $L^{2}(\gamma;\mathcal{H})$. This is provided by the following assumption. {assumption}[Separability Conditions for $L^{2}(\gamma;\mathcal{H})$] We assume that $\gamma$ is a Borel probability measure defined on $\mathcal{B}\left(\Xi\right)$ of a Polish space $\Xi$ and $\mathcal{H}$ is a separable Hilbert space. Under Appendix A, the separability of $L^{2}(\gamma;\mathcal{H})$ follows from a sequence of classical results. Since $\gamma$ is a Borel probability measure on the Polish space $\Xi$, the space $L^{2}(\gamma)$ is a separable Hilbert space (see Proposition 3.4.5 in [8]). By assumption, $\mathcal{H}$ is also separable Hilbert space. Therefore, both $\mathcal{H}$ and $L^{2}(\gamma)$ admit countable orthonormal bases, as guaranteed by Theorem 9 in [25]. Moreover, the algebraic tensor product $L^{2}(\gamma)\otimes\mathcal{H}$ is dense in the Bochner space $L^{2}(\gamma;\mathcal{H})$ ( Lemma 1.2.19 of [47]). Combining these points yields the separability of $L^{2}(\gamma;\mathcal{H})$.
The algebraic tensor product used above is defined by
$f\otimes g(\xi,x)=f(\xi)g(x),\quad\text{ for }f\in L^{2}(\gamma),\,g\in\mathcal{H},$
and finite linear combinations of such elements form $L^{2}(\gamma)\otimes\mathcal{H}$. The inner product of $L^{2}(\gamma;H)$ respects the tensor structure:
[…]
Since $L^{2}(\gamma;H)$ is separable, it admits a countable orthonormal basis $\left\{\tilde{\phi}_{n}\right\}_{n\in\mathbb{N}}\in L^{2}(\gamma;H)$. It is easy to verify that any $q\in L^{2}(\gamma;\mathcal{H})$ can be represented either as
(34) $q(x;\xi)=\sum_{n\in\mathbb{N}}a_{n}(\xi)\,\psi_{n}(x),\quad\text{ for some }\left\{a_{n}\right\}_{n\in\mathbb{N}}\subset L^{2}(\gamma),$
or
(35) $q(x;\xi)=\sum_{n\in\mathbb{N}}b_{n}(x)\,\phi_{n}(\xi),\quad\text{ for some }\left\{b_{n}\right\}_{n\in\mathbb{N}}\subset H.$
where $\left\{\psi_{n}\right\}_{n\in\mathbb{N}}$ and $\left\{\phi_{n}\right\}_{n\in\mathbb{N}}$ are orthonormal bases of $\mathcal{H}$ and $L^{2}(\gamma)$, respectively. Both representations are understood in the $L^{2}(\gamma;\mathcal{H})$ sense.
Theorem A.2 (Universal Approximation via PCE-NN in Neural Measure Spaces).
Suppose the Hilbert space $\mathcal{H}$ satisfies Appendix A for some $k,\,\beta,\,U\subset\mathbb{R}^{in}$, and consider representation of $L^{2}(\gamma;\mathcal{H})$ given in Eq. 35. Let $\rho\in\overline{C_{b}^{k}(\mathbb{R})}^{\beta}$ to be a non polynomial activation function and define the class of neural networks $\mathcal{N}_{\rho,\mathbb{R}^{d_{\text{in}}},d_{\text{out}}}^{(K)}$ as in Eq. 32. For integers $K,\,M\geq 1$, set
$\mathcal{V}_{\mathcal{N}}^{(K,M)}\coloneqq\left\{\begin{split}g_{\theta}:\mathbb{R}^{d_{\text{in}}}\to\mathbb{R}^{d_{\text{out}}M}:\,\text{ for some }g_{\theta}=\left(g_{\theta}^{(1)},\dots,g_{\theta}^{(M)}\right)&\\ &~\hskip-85.35826pt\text{ with }g_{\theta}^{(1)},\dots,g_{\theta}^{(M)}\in\mathcal{N}_{\rho,\mathbb{R}^{d_{\text{in}}},d_{\text{out}}}^{(K)}\end{split}\right\},$
and define the associated class of Neural Measures by
$\mathcal{V}_{\mathcal{N}}^{(K,M),*}=\left\{\begin{split}\left[X_{\theta}\right]_{\!\#}\gamma\in\mathcal{P}_{2}^{\gamma}\left(\,\,\mathcal{H}\,\,\right):\text{ where }X_{\theta}(\xi)=\sum_{n=1}^{M}g_{\theta}^{(n)}\,&\phi_{n}(\xi)\,\,\\ &\text{ for some }g_{\theta}\in\mathcal{V}_{\mathcal{N}}^{(K,M)}\end{split}\right\}.$
Finally, let
$\hat{\mathcal{V}}_{\mathcal{N}}^{*}=\bigcup_{K,M\in\mathbb{N}}\mathcal{V}_{\mathcal{N}}^{(K,M),*}.$
Then, the closure of $\hat{\mathcal{V}}_{\mathcal{N}}^{*}$ with respect to the $W_{2}$-metric on $\mathcal{P}_{2}^{\gamma}(\mathcal{H})$ coincides with the entire space $\mathcal{P}_{2}^{\gamma}(\mathcal{H})$. In other words, for every $\nu\in\mathcal{P}_{2}^{\gamma}(\mathcal{H})$, there exists a sequence $\left\{\mu_{m}\right\}_{m\in\mathbb{N}}\subset\hat{\mathcal{V}}_{\mathcal{N}}^{*}$ such that
$\lim_{m\to\infty}W_{2}(\nu,\mu_{m})=0.$
Remark A.3.
In Theorem A.2, parameters (K,M)(\!K,\!M) act as hyper-parameters. Moreover, recall that in Section 3.2 we fixed a specific neural network architecture $\mathcal{V}_{\mathcal{N}}$, i.e., we fixed K,MK,\,M in $\mathcal{V}_{\mathcal{N}}^{(K,M)}$. Thus, the class $\mathcal{V}_{\mathcal{N}}^{*}$ defined in Eq. 15 corresponds exactly to the specialised case $\hat{\mathcal{V}}_{\mathcal{N}}^{(K,M),*}$ for some fixed K,MK,\,M, under the notation of this section.
Proof A.4.
For any $\mu,\,\nu\in\mathcal{P}_{2}^{\gamma}(\mathcal{H})$, by definition there exist two functions $g,\,f\in L^{2}\left(\gamma;\mathcal{H}\right)$ such that
$\mu(B)=\left[g\right]_{\!\#}\gamma(B),\quad\nu(B)=\left[f\right]_{\!\#}\gamma(B),\qquad\forall B\in\mathcal{B}\left(\mathcal{H}\right).$
Considering the 2-Wasserstein distance $W_{2}\left(\nu,\mu\right)$ on $\mathcal{H}$ (with the norm-induced metric), we have
(36) $\begin{split}W_{2}\left(\nu,\mu\right)^{2}&=\inf_{\pi\in\Pi\left(\nu,\mu\right)}\int_{\mathcal{H}\times\mathcal{H}}\left\|h_{1}-h_{2}\right\|_{\mathcal{H}}^{2}\,\pi\left({\rm d}h_{1},{\rm d}h_{2}\right)\\ &\leq\int_{\Xi}\left\|f(\xi)-g(\xi)\right\|_{\mathcal{H}}^{2}\,\gamma\left({\rm d}\xi\right)=\left\|f-g\right\|_{L^{2}\left(\gamma;\mathcal{H}\right)}^{2}.\end{split}$
The inequality holds since the joint distribution $\left[\left(f,g\right)\right]_{\!\#}\gamma$ is contained within plans $\Pi\left(\nu,\mu\right)$.
Now fix $\nu\in\mathcal{P}_{2}^{\gamma}(\mathcal{H})$, let $f\in L^{2}\left(\gamma;\mathcal{H}\right)$ be such that $\nu=\left[f\right]_{\!\#}\gamma$. Using the expansion Eq. 35, define truncated approximations
(37) $f^{(m)}(x;\xi)=\sum_{n\leq n_{m}}\tilde{b}_{n}(x)\phi_{n}\left(\xi\right),$
where $\left\{\phi_{n}\right\}_{n\in\mathbb{N}}$ is an orthonormal bases of $L^{2}(\gamma)$, $\left\{\tilde{b}_{n}\right\}_{n\in\mathbb{N}}\subset\mathbb{R}$, and $n_{m}$ is an increasing sequence such that
$\left\|f-\sum_{n\leq n_{m}}\tilde{b}_{n}\phi_{n}\right\|_{L^{2}(\gamma;\mathcal{H})}\leq\frac{1}{2\,m}.$
By Appendices A and A.1, we get the closedness of neural networks, i.e.,
$\operatorname{cl}_{\left\|\cdot\right\|_{\mathcal{H}}}\left(\mathcal{N}_{\rho,\mathbb{R}^{d_{\text{in}}},d_{\text{out}}}\Big|_{U}\right)=\mathcal{H}.$
Thus, we can approximate each coefficient function $\tilde{b}_{n}$ with neural network $g_{n}^{(m)}$ so that
(38) $\max_{n\leq n_{m}}\left\|\tilde{b}_{n}-g_{n}^{(m)}\right\|_{\mathcal{H}}<\frac{1}{2\,n_{m}\,m},$
and define
$X_{m}(x;\xi)=\sum_{n\leq n_{m}}g_{n}^{(m)}\,\phi_{n}(\xi),\quad\mu_{m}\coloneqq\left[X_{m}\right]_{\!\#}\gamma.$
Then, the $\left\{\mu_{m}\right\}_{m\in\mathbb{N}}$ sequence satisfies
(39) $\begin{split}W_{2}\left(\nu,\mu_{m}\right)&\leq\left\|f-X_{m}\right\|_{L^{2}\left(\gamma;\mathcal{H}\right)}\\ &\leq\left\|f-f^{(m)}\right\|_{L^{2}\left(\gamma;\mathcal{H}\right)}+\left\|f^{(m)}-X_{m}\right\|_{L^{2}\left(\gamma;\mathcal{H}\right)}\leq\frac{1}{m},\end{split}$
where in the last inequality, we use Eq. 38 and the triangle inequality.
In the case of Galerkin-NN setting, we shall verify below the density on $\mathcal{P}_{2}^{\gamma}(\mathcal{H}_{h})$, where $\mathcal{H}_{h}$ is a Galerkin type subspace of $\mathcal{H}$. The density on $\mathcal{P}_{2}^{\gamma}(\mathcal{H})$ will depend on the approximation Properties of $\mathcal{H}_{h}$ and once these are given it will follow by similar arguments as in the proof of Theorem A.2.
To this end, $\mathcal{H}_{h}\subset\mathcal{H}$ be a fixed finite dimensional Galerkin subspace of dimension $M$, with basis $\left\{\psi_{n}\right\}_{n=1}^{M}$. Suppose a probability space $\left(\mathbb{R}^{d_{\text{ran}}},\mathcal{B}\left(\mathbb{R}^{d_{\text{ran}}}\right),\gamma\right)$ is given such that $L^{2}(\gamma)$ satisfies Appendix A for some hyper-parameters [^2] Example 2.6 of [34] specifies suitable hyperparameters and properties of $\gamma$. $k,\,\beta$, and let $U\coloneqq supp\left(\gamma\right)\subset\mathbb{R}^{d_{\text{ran}}}$. Under the representation Eq. 34, the space $L^{2}(\gamma;\mathcal{H}_{h})$ reduces to the finite expansion
(40) $q(x;\xi)=\sum_{n=1}^{M}a_{n}(\xi)\psi_{n}(x),$
where $\psi_{n}$ are not necessarily orthonormal.
Under these conditions, we consider Neural Measure Spaces as follows: Let $\rho\in\overline{C_{b}^{k}(\mathbb{R})}^{\beta}$ be a non-polynomial activation function. We define class of Neural Measures with finite $K\geq 1$ as
$\mathcal{V}_{\mathcal{N}}^{(K),*}=\left\{\begin{split}\left[X_{\theta}\right]_{\!\#}\gamma\in\mathcal{P}_{2}^{\gamma}\left(\,\,\mathcal{H}_{h}\,\,\right):\text{ where }X_{\theta}(\xi)=\sum_{n=1}^{M}g_{\theta}^{(n)}(\xi)\,\psi_{n}&,\\ &~\hskip-85.35826pt\text{for some }g_{\theta}^{(1)},\dots,g_{\theta}^{(M)}\in\mathcal{N}_{\rho,\mathbb{R}^{d_{\text{ran}}},1}^{(K)}\end{split}\right\}.$
based on the Neural Networks $\mathcal{N}_{\rho,\mathbb{R}^{d_{\text{ran}}},1}^{(M)}$ defined in Eq. 32.
Let us now verify that
$\hat{\mathcal{V}}_{\mathcal{N}}^{*}=\bigcup_{K\in\mathbb{N}}\mathcal{V}_{\mathcal{N}}^{(K),*}$
is dense in $\mathcal{P}_{2}^{\gamma}(\mathcal{H}_{h})$ with respect to the $W_{2}$-metric.
For fixed $\nu\in\mathcal{P}_{2}^{\gamma}(\mathcal{H}_{h})$, let $f\in L^{2}\left(\gamma;\mathcal{H}_{h}\right)$ be such that $\nu=\left[f\right]_{\!\#}\gamma$. According to expansion Eq. 40, we define the coefficient function $\tilde{a}_{n}$ of $f$ by
$f(x;\xi)=\sum_{n=1}^{M}\tilde{a}_{n}(\xi)\psi_{n}\left(x\right).$
Since $L^{2}(\gamma)$ satisfies Appendices A and A.1 provides the following closedness result:
$\operatorname{cl}_{\left\|\cdot\right\|_{L^{2}(\gamma)}}\left(\mathcal{N}_{\rho,\mathbb{R}^{d_{\text{ran}}},1}\Big|_{U}\right)=L^{2}(\gamma).$
Therefore, there exists a sequence of $\left\{g_{1}^{(m)},\dots g_{M}^{(m)}\right\}_{m=1}^{\infty}$ with $g_{n}^{(m)}\in\mathcal{N}_{\rho,\mathbb{R}^{d_{\text{ran}}},1}$ such that
(41) $\max_{n\leq M}\left\|\tilde{a}_{n}-g_{n}^{(m)}\right\|_{L^{2}(\gamma)}<\frac{1}{m\sum_{n=1}^{M}\|\psi_{n}\|_{\mathcal{H}}}.$
Based on the sequence of $g_{n}^{(m)}$’s, we define
$X_{m}(x;\xi)=\sum_{n=1}^{M}g_{n}^{(m)}(\xi)\,\psi_{n}(x),\quad\text{ and set }\quad\mu_{m}\coloneqq\left[X_{m}\right]_{\!\#}\gamma.$
By applying Inequality Eq. 36, we have
$W_{2}\left(\nu,\mu_{m}\right)\leq\left\|f-X_{m}\right\|_{L^{2}\left(\gamma;\mathcal{H}\right)}\leq\max_{n\leq M}\left\|\tilde{a}_{n}-g_{n}^{(m)}\right\|_{L^{2}(\gamma)}\,\,\sum_{n=1}^{M}\|\psi_{n}\|_{\mathcal{H}}<\frac{1}{m}.$
confirming the convergence $\mu_{m}\to\nu$ with respect to $W_{2}$-metric.
UAT for Fully Network-Based Neural Measure Spaces
In the setting of Fully Network-Based Neural Measure Spaces, we now consider a Hilbert space $\mathcal{H}$ tied to our original problem. Specifically, we take $\mathcal{H}=W^{k,2}(D;\mathbb{R}^{d_{\text{out}}})$, the Sobolov space of $\mathbb{R}^{d_{\text{out}}}$-valued functions on a bounded open domain $D\subset\mathbb{R}^{d_{\text{in}}}$, defined as:
$W^{k,2}(D;\mathbb{R}^{d_{\text{out}}})\coloneqq\left\{f\in L^{2}\left(D;\mathbb{R}^{d_{\text{out}}}\right):\,\,\left\|f\right\|_{k,2}^{2}=\sum_{|a|\leq k}\left\|\partial_{a}f\right\|_{L^{2}\left(D;\mathbb{R}^{d_{\text{out}}}\right)}^{2}<\infty\right\},$
where $\partial_{a}f$ denotes the weak derivative of $f$ on some multi-index $a$ (as defined in equation Eq. 31). Since $L^{2}\left(D;\mathbb{R}^{d_{\text{out}}}\right)$ is a separable Hilbert space, it follows that $W^{k,2}(D;\mathbb{R}^{d_{\text{out}}})$ is also separable.
Let $\gamma$ be a Borel probability measure on $\mathcal{B}\left(\Xi\right)$ with $\Xi\subset\mathbb{R}^{d_{\text{ran}}}$. Define the probability measure $\lambda$ on $\mathbb{R}^{d_{\text{in}}}$, as $\lambda({\rm d}x)=\frac{1}{\left|D\right|}1(x\in D){\rm d}x$ and consider the product measure $\gamma\otimes\lambda$ on $\Xi\times D$.
Now define the function space
[…]
i.e., we only differentiate with respect to the input variables $x\in D$.
By applying Fubini’s theorem on functions $f$ which either in $L^{2}\!\!\left(\!\gamma;\!W^{k,2}\!\left(\!D;\mathbb{R}^{d_{\text{out}}\!}\right)\right)$, or in $\Lambda^{(0,k),2}(\gamma\otimes\lambda;\mathbb{R}^{d_{\text{out}}})$, we obtain
$\left\|f\right\|_{L^{2}(\gamma;W^{k,2}(D;\mathbb{R}^{d_{\text{out}}}))}=\left|D\right|\left\|f\right\|_{\Lambda^{(0,k),2}(\gamma\otimes\lambda;\mathbb{R}^{d_{\text{out}}})}.$
From which, we get that these two spaces are isometrically isomorphic, in our case these spaces are identical.
Under an appropriate choice of the hyperparameters $k,\,\beta,\,\Xi\times D\subset\mathbb{R}^{d_{\text{in}}+d_{\text{ran}}},\,\gamma$, one can ensure that $\Lambda^{(0,k),2}(\gamma\otimes\lambda;\mathbb{R}^{d_{\text{out}}})$ satisfies Appendix A. Suitable choices of these hyperparameters can be obtained by combining the conditions listed in Example 2.6 of [34].
Now let $\rho\in\overline{C_{b}^{k}(\mathbb{R})}^{\beta}$ to be a non-polynomial activation function, and define the class of neural networks $\mathcal{N}_{\rho,\mathbb{R}^{d_{\text{in}}+d_{\text{ran}}},d_{\text{out}}}^{(M)}$ as in Eq. 32. By Theorem A.1, we then obtain the approximation result:
$\operatorname{cl}{\left\|\cdot\right\|_{X}}\left(\mathcal{N}_{\rho,\mathbb{R}^{d_{\text{in}}+d_{\text{ran}}},d_{\text{out}}}\Big|_{\Xi\times D}\right)=L^{2}(\gamma;W^{k,2}(D;\mathbb{R}^{d_{\text{out}}})),$
where norm $\left\|\cdot\right\|_{X}$ stands for $L^{2}(\gamma;W^{k,2}(D;\mathbb{R}^{d_{\text{out}}}))$-norm.
Finally, similar to the proof of Theorem A.2 – and more precisely, by applying inequality Eq. 39 which shows that the $L^{2}$-norm controls the $W_{2}$-metric – we obtain the Universal Approximation Theorem for Fully Network-Based Neural Measure Spaces when the target Hilbert space is a Sobolev space. That is, for every $\nu\in\mathcal{P}_{2}^{\gamma}(\mathcal{H})$, there exists a sequence $\left\{\mu_{m}\right\}_{m\in\mathbb{N}}\subset\hat{\mathcal{V}}_{\mathcal{N}}^{*}$ such that
$\lim_{m\to\infty}W_{2}(\nu,\mu_{m})=0.$
A.1 Well-Definedness of $U(\xi)$ Distribution
Based on the Random PDE stated in Equation 6, we are interested in approximating its probability distribution, namely $\left[U\right]_{\!\#}\gamma$. We would like to know under what assumptions can we ensure that this distribution is well-defined. Moreover, does $\left[U\right]_{\!\#}\gamma$ depend on the choice of the underlying probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$?
To answer this, we first define two mappings. Given a stochastic process
$f:D\times\Xi\to\mathbb{R}\quad\text{ with }\quad f(\cdot;\xi)\in\tilde{\mathcal{H}}$,
and
$g:\Gamma\times\Xi\to\mathbb{R}\quad\text{ with }\quad g(\cdot;\xi)\in\overline{\mathcal{H}},$
we define the evaluation map
$Q_{f,g}:\Xi\to\tilde{\mathcal{H}}\times\overline{\mathcal{H}}\times\Xi\quad\text{ with }\quad\xi\mapsto\left(f(\cdot;\xi),g(\cdot;\xi),\xi\right).$
Next, we introduce the solution operator
[…]
where $\tilde{u}(\cdot;\xi)$ is the solution to the Random PDE (similar to the one in Equation 6) for a fixed $\xi=\tilde{\xi}$, and with right hand side replaced by $\tilde{f}$ and $\tilde{g}$.
Then, the solution $U(\xi)$ to the original problem can be represented as:
$U(\xi)=G(f(\cdot;\xi),g(\cdot;\xi),\xi)=G\circ Q_{f,g}(\xi).$
A sufficient condition for ensuring both the well-definedness of $\left[U\right]_{\!\#}\gamma$ and its invariance with respect to the abstract probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ is the following:
  • The solution operator $G$ is well-defined for every point in $\tilde{\mathcal{H}}\times\overline{\mathcal{H}}\times\Xi$, and
  • The composition $G\circ Q_{f,g}(\xi)$ is measurable on each instance of $\left(f,g\right)\in L^{2}(\gamma;\tilde{\mathcal{H}})\times L^{2}(\gamma;\overline{\mathcal{H}})$.
[…]
$\left[U\circ\xi\right]_{\!\#}\mathbb{P}=\left[G\circ Q_{f,g}\circ\xi\right]_{\!\#}\mathbb{P}=\left[G\circ Q_{f,g}\right]_{\!\#}\gamma=\left[G\circ Q_{f,g}\circ\tilde{\xi}\right]_{\!\#}\tilde{\mathbb{P}}=\left[U\circ\tilde{\xi}\right$$_{\!\#}\mathbb{P}$.
This confirms that the push-forward measure $\left[U\right]_{\!\#}\gamma$ is independent of the choice of the underlying probability space, provided the stated conditions are met.
A.2 Discrete Neural Network Spaces
We describe our notation regarding the generic discrete neural network spaces which we denote by $V_{\mathcal{N}}.$ These are spaces approximating functions
$v:\mathbb{R}^{d_{I}}\to\mathbb{R}^{d_{O}}.$
We have used different choices for $d_{I}$ and $d_{O}$ throughout this paper. To fix ideas, we present a generic structure of the spaces $V_{\mathcal{N}}.$ Of course several alternative architectures are possible. A deep neural network maps every point $\overline{y}\in\mathbb{R}^{d_{I}}$ to $v_{\theta}(\overline{y})\in\mathbb{R}^{d_{O}}$, through
(43) $v_{\theta}(\overline{y})=\mathcal{C}_{L}(\overline{y}):=C_{L}\circ\sigma\circ C_{L-1}\cdots\circ\sigma\circ C_{1}(\overline{y})\quad\forall\overline{y}\in\mathbb{R}^{d_{I}}.$
The intermediate layers $C_{k}$, are affine maps of the form
(44) $C_{k}y=W_{k}y+b_{k},\qquad\text{where }W_{k}\in\mathbb{R}^{d_{k+1}\times d_{k}},b_{k}\in\mathbb{R}^{d_{k+1}},$
where the dimensions $d_{k}$ may vary with each layer $k$ and $\sigma(y)$ denotes the vector with the same number of components as $y$, where $\sigma(y)_{i}=\sigma(y_{i}).$ The index $\theta$ represents collectively all the parameters of the network. The space of parameters is denoted by $\Theta$:
$\Theta=\left\{\theta:=\left(b_{1},W_{1},\dots W_{L},b_{L}\right):\text{ for }b_{i}\in\mathbb{R}^{d_{k}},\,W_{i}\in\mathbb{R}^{d_{k+1}\times d_{k}},\ k=1,\dots,L\right\}.$
The set of networks $\mathcal{C}_{L}$ with a given architecture of the form Eq. 43, Eq. 12 is called $\mathcal{N}$. Then $\Theta\in\mathbb{R}^{\dim{\mathcal{N}}}$ where the total number of degrees of freedom of $\mathcal{N},$ is $\dim{\mathcal{N}}=\sum_{k=1}^{L}d_{k+1}(d_{k}+1).$ The nonlinear discrete set of functions $V_{\mathcal{N}}$ is defined as
(45) $V_{\mathcal{N}}=\{u_{\theta}:\varOmega_{T}\to\mathbb{R},\ \text{where }u_{\theta}(\bm{x})=\mathcal{C}_{L}(\bm{x}),\ \text{for some }\mathcal{C}_{L}\in\mathcal{N}\,\}.$
A.3 Neural Networks with Random Input as Generative Models
In generative modeling, neural networks receive both a standard input and a random input (often called the latent code or noise). Formally, assume a probability distribution $\gamma$ over $\mathbb{R}^{d_{\text{ran}}}$ (also known as the seed or reference distribution), and a neural network $g_{\theta}\in\mathcal{V}_{\mathcal{N}}$. Then, the generative model is the random variable:
$g_{\theta}(x;\xi),\qquad\text{ where }\xi\sim\gamma.$
This induces a probability distribution over outputs, defined via the pushforward measure:
$g_{\theta}(x;\xi)\sim\left[g_{\theta}(x,\cdot)\right]_{\!\#}\gamma.$
To simplify notation, define:
$Q_{\theta,x}(\xi):=g_{\theta}(x,\xi)$
Then, the output distribution for a fixed input $x$ is:
$\mu_{\theta}(x,\,A):=\gamma\circ Q_{\theta,x}^{-1}(A)=\int_{\mathbb{R}^{d_{\text{ran}}}}\mathds{1}_{\{g_{\theta}(x,\xi)\in A\}}\,\gamma({\rm d}\xi),\qquad\forall A\in\mathcal{B}\left(\mathbb{R}^{d_{\text{out}}}\right).$
If the mapping $g_{\theta}:\mathbb{R}^{d_{\text{in}}}\times\mathbb{R}^{d_{\text{ran}}}\to\mathbb{R}^{d_{\text{out}}}$ is measurable with respect to the Borel $\sigma$-algebras $\mathcal{B}\left(\mathbb{R}^{d_{\text{in}}}\times\mathbb{R}^{d_{\text{ran}}}\right)$ and $\mathcal{B}\left(\mathbb{R}^{d_{\text{out}}}\right)$, then the following properties hold:
  • For every fixed $x\in\mathbb{R}^{d_{\text{in}}}$, $\mu_{\theta}(x,\cdot)$ is a probability measure on $\mathbb{R}^{d_{\text{out}}}$
  • For every fixed $A\in\mathcal{B}\left(\mathbb{R}^{d_{\text{out}}}\right)$, the function $x\mapsto\mu_{\theta}(x,\,A)$ is measurable with respect to $\mathcal{B}\left(\mathbb{R}^{d_{\text{in}}}\right)$.
  • $\mu_{\theta}(\cdot,A)$ is $\mathcal{B}\left(\mathbb{R}^{d_{\text{in}}}\right)/\mathcal{B}\left(\mathbb{R}^{d_{\text{out}}}\right)$-measurable
Therefore, $g_{\theta}$ defines a stochastic kernel $\mu_{\theta}(x,\,{\rm d}y)$ from $\mathbb{R}^{d_{\text{in}}}$ to $\mathbb{R}^{d_{\text{out}}}$.
To establish these properties, one can utilise the fact that
$\mathcal{B}\left(\mathbb{R}^{d_{\text{in}}}\times\mathbb{R}^{d_{\text{ran}}}\right)=\mathcal{B}\left(\mathbb{R}^{d_{\text{in}}}\right)\otimes\mathcal{B}\left(\mathbb{R}^{d_{\text{ran}}}\right),$
since both $\mathbb{R}^{d_{\text{in}}}$ and $\mathbb{R}^{d_{\text{ran}}}$ are Polish spaces, see Theorem 17.25 in [20].
Figure 4: Training and testing loss over time during the training of the PINN architecture for the bistable ODE Eq. 19.
Figure 5: Training and testing loss function over time of the training of the PINN and the PINN-PCE architecture for the diffusion equation Eq. 21.
Figure 6: Training and testing loss function over time of the training of the PINN and the PINN-PCE architecture for the reaction-diffusion equation Eq. 27.
Acknowledgments
The authors used AI-assisted tools (ChatGPT by OpenAI) for minor editorial assistance, including grammar correction and language polishing. All substantive ideas, analysis, and results presented in this work are entirely the authors’ own.